OpenStudy (anonymous):

Due to plate tectonics, the summit of Mount Everest moves about 4.5x10^-3 meters northeastward in one year. How many more meters does the summit of Mt.Everest move in 11 years? Express your answer in scientific notation.

OpenStudy (anonymous):

462 is the amount of meters that summits Mt. E.

OpenStudy (raden):

4.5x10^-3 x 11 = 49.5 x 10^-3 now, can you evaluate this : 49.5 x 10^-3 - 4.5 x 10^-3 = .... ?

OpenStudy (anonymous):

4.5?

OpenStudy (anonymous):

no the answer to his equation was my answer. I'm sorry i didnt explain it 49.5 x 10^-3 - 4.5 x 10^-3 = 462

OpenStudy (anonymous):

in other way \[\huge\frac{ 49.5 }{ 10^3 } - \frac{ 4.5 }{ 10^3 }\] you can move base 10 along with the exponent to the denominator, this would change the negative exponent to positive, to make the problem simpler

OpenStudy (raden):

what is 49.5 - 4.5 ? then times 10^-3

OpenStudy (anonymous):

^